# Smooth Model Predictive Path Integral Control without Smoothing

### v7

@@ Abstract
Index Terms:
Optimization and optimal control, planning under uncertainty, model learning for control, autonomous vehicle navigation, field robots.
@@ III-B2 Smoothing along the “t-axis”
       ${\bf x}_{0}\leftarrow\text{SubscribeState()}$();
       for $k\leftarrow 0$ to $K-1$ do
@@ IV Experiments on an Inverted Pendulum
Figure 2: Average state costs during optimization procedures with different control methods. Each iteration consists of $20$ time steps. Note that we apply logarithmic scaling to the y-axis and that the costs are clamped to a minimum of $0.01$.
@@
The results show that only SMPPI successfully stabilized the pendulum in the upright position in all cases. As seen in our video, the other methods showed similar results. They continually failed to push the pendulum to the upright position because they did not apply enough torque during the swinging motions. In the case of MPPI baseline i), it was vulnerable to chattering and could not converge into an optimal control sequence when the optimal distribution changed rapidly. Imposing an additional action cost ii), which was employed to alleviate chattering, also failed because this strategy violates the information theoretic interpretation of MPPI while using non-affine dynamics. Methods using external smoothing algorithms iii)--iv) failed due to the negative effects of external smoothing. We discuss these issues further in Section V-D.
@@ V-B Training the Neural Network Vehicle Model
 Input $(\left|{\bf x}\right|+\left|{\bf a}\right|)\times h$
 Hidden Layer 1 2×2$\times$2 ×Input Size ReLU
 Hidden Layer 2 4×4$\times$4 ×Input Size ReLU
 Hidden Layer 3 6×6$\times$6 ×Input Size ReLU
 Hidden Layer 4 2×2$\times$2 ×Input Size Linear
 Output $\left|{\bf x}\right|$
@@
    Zig-zag driving at low speeds ($20-25$ km/h) on the race track.

### v8

@@ III-B2 Smoothing along the “t-axis”
 ${\bf x}_{0}\leftarrow\text{SubscribeState()}$;
 for $k\leftarrow 0$ to K1K-1 do
@@ IV Experiments on an Inverted Pendulum
Figure 2: Average state costs during optimization procedures with different control methods. Each iteration consists of 2020 time steps. Note that we apply logarithmic scaling to the y-axis and that the costs are clamped to a minimum of $0.01$.
@@
The results show that only SMPPI successfully stabilized the pendulum in the upright position in all cases. As seen in our video, the other methods showed similar results. They continually failed to push the pendulum to the upright position because they did not apply enough torque during the swinging motions. In the case of MPPI baseline i), it was vulnerable to chattering and could not converge into an optimal control sequence when the optimal distribution changed rapidly. Imposing an additional action cost ii), which was employed to alleviate chattering, also failed because this strategy violates the information theoretic interpretation of MPPI while using non-affine dynamics. Methods using external smoothing algorithms iii)-iv) failed due to the negative effects of external smoothing. We discuss these issues further in Section V-D.
@@ V-B Training the Neural Network Vehicle Model
 Input $(\left|{\bf x}\right|+\left|{\bf a}\right|)\times h$ -
 Hidden Layer 1 $2\timesInput$ Size ReLU
 Hidden Layer 2 $4\timesInput$ Size ReLU
 Hidden Layer 3 $6\timesInput$ Size ReLU
 Hidden Layer 4 $2\timesInput$ Size Linear
 Output $\left|{\bf x}\right|$ -
@@
    Zig-zag driving at low speeds (202520-25 km/h) on the race track.